# Disentangling data distribution for Federated Learning

### v1

@@ Abstract
Federated Learning (FL) facilitates collaborative training of a global model whose performance is boosted by private data owned by distributed clients, without compromising data privacy. Yet the wide applicability of FL is hindered by entanglement of data distributions across different clients. This paper demonstrates for the first time that by disentangling data distributions FL can in principle achieve efficiencies comparable to those of distributed systems, requiring only one round of communication. To this end, we propose a novel FedDistr algorithm, which employs stable diffusion models to decouple and recover data distributions. Empirical results on the CIFAR100 and DomainNet datasets show that FedDistr significantly enhances model utility and efficiency in both disentangled and near-disentangled scenarios while ensuring privacy, outperforming traditional federated learning methods.
@@ 1 Introduction
Figure 1: Disentangled and near-disentangled cases: In the Disentangled case, two clients have data distributions on two disentangled base distribution, $P_{1}$ and $P_{2}$, separately. In the $\xi$-entangled case, each client has data distributions across both disentangled base distribution $P_{1}$ and $P_{2}$, but with one base distribution dominating the other. In both case, client $k$ ($k=1,2$) communicates with the server through data distribution in a single round (upload the distribution $[[S_{k}]]$ that applying the privacy preserving mechanishm on $S_{k}$). The server employs different aggregation strategies for the two scenarios: $\text{Aggre}_{A}$ and $\text{Aggre}_{B}$ for disentangled and near-disentangled data distributions respectively.
@@
Despite the extensive research in Federated Learning (FL), its practical application remains limited. A key challenge is to achieve high efficiency while preserving both model utility and privacy (McMahan et al., 2017; Kairouz et al., 2021). There is a consensus that this inefficiency stems from the entanglement of data distribution across clients, where many rounds of communications are required to ensure the convergence of the global model (Zhao et al., 2018; X. Li et al., n.d.; Tian et al., 2024). The opposite of entanglement implies that client data distributions at different dimensions are disentangled as shown in Figure 1(a).
@@
We believe that the ideal federated learning algorithm can achieve efficiency levels comparable to those of ideal distributed systems that attain full parallelism (Sunderam, 1990), provided that the data distributions across clients can be entirely disentangled, as illustrated in Fig. 1(a) (see details in Def. 2). As shown by Theorem 1, it is actually a sufficient condition for achieving ideal efficiency of federated learning with only one round of communication being needed. Moreover, this ideal condition is often approximately fulfilled in practice e.g., when millions of mobile device clients participate in federated learning and most client data distributions are different from each other (Sattler et al., 2019; Tian et al., 2024). In other words, under such a near-disentangled condition (see Def. 2 and Fig. 1(b)), there exists a federated learning algorithm capable of achieving global model utility with only one communication round, while ensuring that the utility loss remains within a tolerable threshold (see Theorem 2).
@@
In order to take full advantage of the described disentangled and near–disentangled cases (see Fig. 1), we propose to first disentangle data distributions into distinct components such that each client can launch their respective learning task independently. This disentanglement will lead to significant reduction of required communication rounds as revealed by Theorem 1. Technology-wise, there is a large variety of methods that can be utilized to decompose data distribution, ranging from subspace decomposition (Abdi and Williams, 2010; Von Luxburg, 2007) to dictionary learning (Tošić and Frossard, 2011) etc. (see Section 2.2 for detailed review). In this work, we propose an algorithm, called FedDistr, which leverages the stable diffusion model technique (Croitoru et al., 2023) due to its robust ability to extract and generate data distributions effectively. In our approach, clients locally extract data distributions via stable diffusion model, and then upload these decoupled distributions to the server. The server actively identifies the orthogonal or parallel between the base distributions uploaded by clients and aggregate the orthogonal distribution once.
@@
Previous work (Zhang et al., 2022) has shown that it is impossible to achieve the optimal results among the utility, privacy and efficiency. The proposed FedDistr offers a superior balance between model utility and efficiency under the disentangled and near-disentangled scenarios, while still ensuring the privacy-preserving federated learning: (1) By decoupling client data distributions into the different base distributions, the server actively aligns the base distribution across different client, while FedAvg does not perform this decoupling, leading to a decline in global performance, especially for the disentangled case; (2) The proposed FedDistr requires only one round of communication, and the amount of transmitted distribution parameters is much smaller than that of model gradients; (3) The proposed FedDistr transmits a minimal amount of data distribution parameters, thereby mitigating the risk of individual data privacy leakage to some extent. Furthermore, privacy mechanisms such as differential privacy (DP) can be integrated into FedDistr, offering additional protection for data privacy.
@@ 2.1 Communication Efficient Federated Learning
To address the Non-IID problem, numerous methods have been proposed (T. Li et al., 2020; Arivazhagan et al., 2019), which introduce constraints on local training to ensure that models trained on heterogeneous data do not diverge excessively from the global model. However, most of these methods require numerous communication rounds, which is impractical, particularly in wide-area network settings.
Furthermore, some methods propose a one-shot federated learning approach (Guha et al., 2019; Q. Li et al., 2020), which requires only a single communication round by leveraging techniques such as knowledge distillation or consistent voting. However, significant challenges arise in this context under Non-IID settings (Diao et al., 2023).
@@ 2.2 Distribution Generation
This paper focuses on transferring distribution instead of models. There are several distribution generation methods: Parametric Methods (Reynolds et al., 2009) assume that the data follows a specific distribution with parameters that can be estimated from the data. Gaussian Mixture Model (GMM) represents a distribution as a weighted sum of multiple Gaussian components. Generative Models (Goodfellow et al., 2020; Croitoru et al., 2023) aim to learn the underlying distribution of the data so that new samples can be generated from the learned distribution. For example, GANs Goodfellow et al. (2020); Karras et al. (2020) consist of two networks, a generator $G$ and a discriminator $D$. The generator learns to produce data similar to the training data, while the discriminator tries to distinguish between real and generated data. Principal Component Analysis (PCA) (Abdi and Williams, 2010) is a linear method for reducing dimensionality and capturing the directions of maximum variance. It decomposes the data covariance matrix $\Sigma$ into eigenvectors and eigenvalues. Dictionary Learning and Sparse Coding (Tošić and Frossard, 2011) seeks to represent data as a sparse linear combination of basis vectors (dictionary atoms).
@@ 3.1 Problem Formulation
Numerous studies (Zhao et al., 2018; X. Li et al., n.d.; Tian et al., 2024) have demonstrated that Eq. (1) converges to Eq. (2) under IID settings. However, significant discrepancies arise in Non-IID scenarios, leading to an increased number of communication rounds and a decline in model utility for Eq. (1). While data heterogeneity in extreme Non-IID settings was considered the main cause for FL algorithms having to compromise model performances for reduced rounds of communication, this paper discloses that direct optimization of Eq. (2) can actually be achieved at the cost of a single round of communication. Both theoretical analysis and empirical studies show that data heterogeneity is a blessing rather than a curse, as long as data distributions among different clients can be completely disentangled (see Theorem 1). Moreover, we aim to learn the distribution $S$ in a single communication by transferring either the distribution or its parameters, thereby mitigating the risk of individual data privacy leakage (Xiao and Devadas, 2023) as follows:
$\min_{S}\sum_{k=1}^{K}\lambda_{k}\tilde{f}(S;\mathcal{D}_{k}),\quad\text{s.t. }T_{c}=1$ (3)
@@ 3.2.1 $\xi$-entangled
Assume the data $\mathcal{D}$ is drawn from $m$ distinct base distributions. Therefore, we can decompose the complex distribution $S$ into simpler components, each representing a portion of the overall distribution (Reynolds et al., 2009; Abdi and Williams, 2010):
@@
It is important to note that Entangled coefficient is closely related to the degree of Non-IID data in federated learning (FL) (X. Li et al., n.d.). Specifically, under IID conditions, where each client follows the same independent and identical distribution, this implies that $\pi_{i,k_{1}}=\pi_{i,k_{2}}>0$ for any $k_{1},k_{2}$ in Eq. (6). Consequently, $\xi$ equals $1$. In contrast, under extreme Non-IID conditions, where each client has entirely distinct data distributions (e.g., client $k_{1}$ possesses a dataset consisting solely of ’cat’ images, whereas client $k_{2}$ possesses a dataset exclusively of ’dog’ images), the Entanglement Coefficient, $\xi$, equates to zero.
@@ 3.2.2 Analysis on Disentangled and Near-disentangled case
Numerous studies, as referenced in the literature (Kairouz et al., 2021; X. Li et al., n.d.), have demonstrated that the performance of the Federated Averaging (FedAvg) algorithm (McMahan et al., 2017) is significantly influenced under disentangled conditions. It will be demonstrated herein that there exists an algorithm capable of preserving model utility and efficiency within the disentangled scenario, with the distinct advantage of necessitating merely a single communication round.
@@
$Pr(\epsilon_{u}\leq\epsilon)\geq 1-\exp(-\frac{\min\{n_{1},\cdots,n_{K}\}\epsilon^{2})}{2L^{2}}$ (8)
@@
When the distributions of across $K$ clients satisfy near-disentangled condition, specifically, $\xi<\frac{1}{(K-1)^{2}}$, if $f$ is $L$-lipschitz, then there exists a privacy-preserving federated algorithm that requires only one communication round and achieves a utility loss of less than $\epsilon$ with a probability of at least $1-\exp(-\frac{(1-(K-1)\sqrt{\xi})n\epsilon^{2}}{2mL^{2}})$, i.e.,
$Pr(\epsilon_{u}\leq\epsilon)\geq 1-\exp(-\frac{(1-(K-1)\sqrt{\xi})n\epsilon^{2}}{2mL^{2}})$ (9)
@@ 4 The Proposed Algorithm
This section introduces the proposed algorithm, called FedDistr, achieving one communication round while maintaining the model utility by leveraging the Latent Diffusion Model (LDM) Ho and Salimans (2022) due to its fast inference speed and the wide availability of open-source model parameters. FedDistr is mainly divided into the following three steps (see Fig. 3 and Algo. 1):
@@ 4.1 Distribution Disentangling
The first step is to disentangle the local data distribution into several base distribution for each client by leveraging the autoencoder part of the LDM. The autoencoder Van Den Oord et al. (2017); Agustsson et al. (2017) consists of an encoder, $E$, and a decoder, $\mathcal{D}$. The encoder, $E$, maps input images $x$ to a spatial latent code, denoted as $z=E(x)$, while the decoder $\mathcal{D}$ is trained to perform the inverse mapping of the encoder, achieving an approximation $x\approx\mathcal{D}(E(x))$.
@@ 4.2 Distribution Generation
To estimate the base distribution $P$, each client learns their distribution parameters $v(P)$ by leveraging the embedding inversion technique in the diffusion model of LDM Ho and Salimans (2022); Liang et al. (2024). Specifically, a fixed prompt, such as “a photo of”, is encoded using a frozen text encoder into a word embedding $p$. A learnable distribution parameter (prompt embedding) $v$ is then randomly initialized and concatenated with $p$ to form the guiding condition [$p$; $v$]. This combined condition is employed to compute the LDM’s loss function. Our optimization goal to learn the distribution parameter $v_{k,i}$ of $i_{th}$ base distribution for client $k$ is formulated as:
@@
where $z$ is the latent code of the input image $B_{k,i}$ generated by the encoder $E$, $\epsilon$ refers to noise sampled from the standard normal distribution $\mathcal{N}(0,1)$, $t$ denotes the timestep in the diffusion process, $\alpha_{t}$ is a hyperparameter related to $t$ and $\epsilon_{\theta}$ is the model that predicts the noise. Therefore, we derive the $m_{k}$ data representations $v_{k,i}$ by solving the optimization problem in Eq. (12). These representations are then fed into the diffusion model to generate the corresponding datasets for each disentangled base distribution $i$. In the federated learning setting, each client uploads their respective set of representations ${v_{k,i}}_{i=1}^{m_{k}}$ and corresponding data number to the server.
4.3 Active Distribution Alignment
Moreover, since the server does not have access to the sequence of the distribution parameters, matching these parameters across different clients presents a challenge. This matching problem can be framed as an assignment problem in a bipartite graph Dulmage and Mendelsohn (1958). To address this assignment problem, we utilize the Kuhn-Munkres algorithm (KM) Zhu et al. (2011), which is designed to find the maximum-weight matching in a weighted bipartite graph, or equivalently, to minimize the assignment cost. Specifically, the goal is to optimize the following expression for matching prompt embeddings between client $1$ and $k$:
$\min\sum_{i_{1}=1}^{m_{k_{1}}}\sum_{i_{2}=1}^{m_{k_{2}}}\|v_{k_{1},i_{1}}-v_{k_{2},i_{2}}\|e^{k_{1},k_{2}}_{i_{1},i_{2}}$ (13)
@@ 4.3 Active Distribution Alignment
6:     $z_{k,i}=E(x_{k,i})$
7:     Clustering $\mathcal{D}_{k}$ into $m_{k}$ datasets $\{B_{k,i}\}_{i=1}^{m_{k}}\}$ according to $z_{k,i}$ as Eq. (11).
8:     Initialize $\{v_{k,i}\}_{i=1}^{m_{k}}$.
9:     for $i$ in [$m_{k}$do:
10:         Learn the base distribution parameter $v_{k,i}$ by optimizing Eq. (12)
11:     Upload $\{v_{k,i}\}_{i=1}^{m_{k}}$ to the server;
12$\triangleright$ The server performs:
13:Build the binary assignment $e_{i,j}^{k_{1},k_{2}}$ via Eq. (13)
14:Obtain multiple parallel set $\mathcal{I}_{p}=\{(k,i)|e^{k_{1},k_{2}}_{i_{1},i_{2}}=1,k\in[K],i\in[m_{k}]\}$
15:Obtain the orthogonal set $\mathcal{I}_{o}=\{(k,i)|k\in[K],i\in[m_{k}]\}-\mathcal{I}_{p}$
16:Obtain $v_{o}=\bigcup_{(k,i)\in\mathcal{I}_{o}}v_{k,i}$
17:Select the distribution parameter $v_{p}$ in $\mathcal{I}_{p}$ with the maximum trained data number;
18:Distribute $v=[v_{p},v_{o}]$ to all clients;
19:return $v$

### v2

@@ Abstract
Federated Learning (FL) facilitates collaborative training of a global model whose performance is boosted by private data owned by distributed clients, without compromising data privacy. Yet the wide applicability of FL is hindered by the entanglement of data distributions across different clients. This paper demonstrates for the first time that by disentangling data distributions FL can in principle achieve efficiencies comparable to those of distributed systems, requiring only one round of communication. To this end, we propose a novel FedDistr algorithm, which employs stable diffusion models to decouple and recover data distributions. Empirical results on the CIFAR100 and DomainNet datasets show that FedDistr significantly enhances model utility and efficiency in both disentangled and near-disentangled scenarios while ensuring privacy, outperforming traditional federated learning methods.
@@ 1 Introduction
Figure 1: Disentangled and near-disentangled cases: In the Disentangled case, two clients have data distributions on two disentangled base distribution, $P_{1}$ and $P_{2}$, separately. In the $\xi$-entangled case, each client has data distributions across both disentangled base distribution $P_{1}$ and base distribution $P_{2}$, but with one base distribution dominating the other. In both case, client $k$ ($k=1,2$) communicates with the server through data distribution in a single round (upload the distribution $[[S_{k}]]$ that applying the privacy preserving mechanishm on $S_{k}$). The server employs different aggregation strategies for the two scenarios: $\text{Aggre}_{A}$ and $\text{Aggre}_{B}$ for disentangled and near-disentangled data distributions respectively.
@@
Despite the extensive research in Federated Learning (FL), its practical application remains limited. A key challenge is to achieve high efficiency while preserving both model utility and privacy (McMahan et al., 2017; Kairouz et al., 2021). There is a “consensus” that this inefficiency stems from the disentanglement of data distribution across clients, where many rounds of communications are required to ensure the convergence of the global model (Zhao et al., 2018; Li et al., ; Tian et al., 2024). In contrast, tasks with tight disentanglement exhibit greater efficiency in distributed systems Coulouris et al. (2005). This is because complex tasks can be broken down into disentangled subtasks, which can be assigned to multiple clients for parallel execution. For example, a medical imaging diagnosis task for HIV and Alzheimer’s disease can be disentangled into two distinct sub-tasks. These tasks are assigned to two hospitals, each training its respective model parallelly and unified to the one global model.
@@
We believe that the ideal federated learning algorithm can achieve efficiency levels comparable to those of ideal distributed systems that attain full parallelism (Sunderam, 1990), provided that the data distributions across clients can be entirely disentangled, as illustrated in Def. 2. As shown by Theorem 1, it is actually a sufficient condition for achieving ideal efficiency of federated learning with only one round of communication being needed. Moreover, this ideal condition is often approximately fulfilled in practice e.g., when millions of mobile device clients participate in federated learning and most client data distributions are different from each other (Sattler et al., 2019; Tian et al., 2024). In other words, under such a near-disentangled condition (see Def. 2), there exists a federated learning algorithm capable of achieving global model utility with only one communication round, while ensuring that the utility loss remains within a tolerable threshold (see Theorem 2).
@@
In order to take full advantage of the described disentangled and near–disentangled cases (see Fig. 1), we propose to first disentangle data distributions into distinct components such that each client can launch their respective learning task independently. This disentanglement will lead to significant reduction of required communication rounds as revealed by Theorem 1. Technology-wise, there is a large variety of methods that can be utilized to decompose data distribution, ranging from subspace decomposition (Abdi & Williams, 2010; Von Luxburg, 2007) to dictionary learning (Tošić & Frossard, 2011) etc. (see Section 2.2 for detailed review). In this work, we propose an algorithm, called FedDistr, which leverages the stable diffusion model technique (Croitoru et al., 2023) due to its robust ability to extract and generate data distributions effectively. In our approach, clients locally extract data distributions via a diffusion model, and then upload these decoupled distributions to the server. The server actively identifies the orthogonal or parallel between the base distributions uploaded by clients and aggregates the orthogonal distribution once.
@@
Previous work (Zhang et al., 2022) has shown that it is impossible to achieve the optimal results among the utility, privacy and efficiency. The proposed FedDistr offers a superior balance between model utility and efficiency under the disentangled and near-disentangled scenarios, while still ensuring the privacy-preserving federated learning: (1) By disentangling client data distributions into the different base distributions, the server actively aligns the base distribution across different client, while FedAvg does not perform this disentangling, leading to a decline in global performance, especially for the disentangled case; (2) The proposed FedDistr requires only one round of communication, and the amount of transmitted distribution parameters is much smaller than that of model gradients; (3) The proposed FedDistr transmits a minimal amount of data distribution parameters, thereby mitigating the risk of individual data privacy leakage to some extent. Furthermore, privacy mechanisms such as differential privacy (DP) can be integrated into FedDistr, offering additional protection for data privacy.
@@ 2.1 Communication Efficient Federated Learning
To address the Non-IID problem, numerous methods have been proposed (Li et al., 2020b; Arivazhagan et al., 2019), which introduce constraints on local training to ensure that models trained on heterogeneous data do not diverge excessively from the global model. However, most of these methods require numerous communication rounds, which is impractical, particularly in wide-area network settings.
Furthermore, some methods propose a one-shot federated learning approach (Guha et al., 2019; Li et al., 2020a), which requires only a single communication round by leveraging techniques such as knowledge distillation or consistent voting. However, significant challenges arise in this context under Non-IID settings (Diao et al., 2023).
@@ 2.2 Distribution Generation
This paper focuses on transferring distribution instead of models. There are several distribution generation methods: Parametric Methods (Reynolds et al., 2009) assume that the data follows a specific distribution with parameters that can be estimated from the data. Gaussian Mixture Model (GMM) represents a distribution as a weighted sum of multiple Gaussian components. Generative Models (Goodfellow et al., 2020; Croitoru et al., 2023) aim to learn the underlying distribution of the data so that new samples can be generated from the learned distribution. For example, GANs Goodfellow et al. (2020); Karras et al. (2020) consist of two networks, a generator $G$ and a discriminator $D$. The generator learns to produce data similar to the training data, while the discriminator tries to distinguish between real and generated data. Principal Component Analysis (PCA) (Abdi & Williams, 2010) is a linear method for reducing dimensionality and capturing the directions of maximum variance. It decomposes the data covariance matrix $\Sigma$ into eigenvectors and eigenvalues. Dictionary Learning and Sparse Coding (Tošić & Frossard, 2011) seeks to represent data as a sparse linear combination of basis vectors (dictionary atoms).
@@ 3.1 Problem Formulation
Numerous studies (Zhao et al., 2018; Li et al., ; Tian et al., 2024) have demonstrated that Eq. (1) converges to Eq. (2) under IID settings. However, significant discrepancies arise in Non-IID scenarios, leading to an increased number of communication rounds and a decline in model utility for Eq. (1). While data heterogeneity in extreme Non-IID settings was considered the main cause for FL algorithms having to compromise model performances for reduced rounds of communication, this paper discloses that direct optimization of Eq. (2) can actually be achieved at the cost of a single round of communication. Both theoretical analysis and empirical studies show that data heterogeneity is a blessing rather than a curse, as long as data distributions among different clients can be completely disentangled (see Theorem 1). Moreover, we aim to learn the distribution $S$ in a single communication by transferring either the distribution or its parameters, thereby mitigating the risk of individual data privacy leakage (Xiao & Devadas, 2023) as follows:
$\min_{S}\sum_{k=1}^{K}\lambda_{k}\tilde{f}(S;\mathcal{D}_{k}),\quad\text{s.t. }T_{c}=1.$ (3)
@@ 3.2.1 $\xi$-entangled
Assume the data $\mathcal{D}$ is drawn from $m$ distinct base distributions. Therefore, we can decompose the complex distribution $S$ into simpler components, each representing a portion of the overall distribution (Reynolds et al., 2009; Abdi & Williams, 2010):
@@
It is important to note that Entangled coefficient is closely related to the degree of Non-IID data in federated learning (FL) (Li et al., ). Specifically, under IID conditions, where each client follows the same independent and identical distribution, this implies that $\pi_{i,k_{1}}=\pi_{i,k_{2}}>0$ for any $k_{1},k_{2}$ in Eq. (6). Consequently, $\xi$ equals $1$. In contrast, under extreme Non-IID conditions, where each client has entirely distinct data distributions (e.g., client $k_{1}$ possesses a dataset consisting solely of ‘cat’ images, whereas client $k_{2}$ possesses a dataset exclusively of ‘dog’ images), the Entanglement Coefficient, $\xi$, equates to zero.
@@ 3.2.2 Analysis on Disentangled and Near-disentangled case
Numerous studies, as referenced in the literature (Kairouz et al., 2021; Li et al., ), have demonstrated that the performance of the Federated Averaging (FedAvg) algorithm (McMahan et al., 2017) is significantly influenced under disentangled conditions. It will be demonstrated herein that there exists an algorithm capable of preserving model utility and efficiency within the disentangled scenario, with the distinct advantage of necessitating merely a single communication round.
@@
$Pr(\epsilon_{u}\leq\epsilon)\geq 1-\exp(-\frac{\min\{n_{1},\cdots,n_{K}\}\epsilon^{2})}{2L^{2}}.$ (8)
@@
When the distributions across $K$ clients satisfy near-disentangled condition, specifically, $\xi<\frac{1}{(K-1)^{2}}$, if $f$ is $L$-lipschitz, then there exists a privacy-preserving federated algorithm that requires only one communication round and achieves a utility loss of less than $\epsilon$ with a probability of at least $1-\exp(-\frac{(1-(K-1)\sqrt{\xi})n\epsilon^{2}}{2mL^{2}})$, i.e.,
$Pr(\epsilon_{u}\leq\epsilon)\geq 1-\exp(-\frac{(1-(K-1)\sqrt{\xi})n\epsilon^{2}}{2mL^{2}}).$ (9)
@@ 4 The Proposed Algorithm
This section introduces the proposed algorithm, called FedDistr, achieving one communication round while maintaining the model utility by leveraging the Latent Diffusion Model (LDM, Ho & Salimans, 2022) due to its fast inference speed and the wide availability of open-source model parameters. FedDistr is mainly divided into the following four steps (see Fig. 3 and Algo. 1):
@@ 4.1 Distribution Disentangling
The first step is to disentangle the local data distribution into several base distribution for each client by leveraging the autoencoder part of the LDM. The autoencoder (Van Den Oord et al., 2017; Agustsson et al., 2017) consists of an encoder, $E$, and a decoder, $\mathcal{D}$. The encoder, $E$, maps input images $x$ to a spatial latent code, denoted as $z=E(x)$, while the decoder $\mathcal{D}$ is trained to perform the inverse mapping of the encoder, achieving an approximation $x\approx\mathcal{D}(E(x))$.
@@ 4.2 Distribution Generation
To estimate the base distribution $P$, each client learns their distribution parameters $v(P)$ by leveraging the embedding inversion technique in the diffusion model of LDM (Ho & Salimans,2022;Liang et al.,2024). Specifically, a fixed prompt, such as “a photo of”, is encoded using a frozen text encoder into a word embedding $p$. A learnable distribution parameter (prompt embedding) $v$ is then randomly initialized and concatenated with $p$ to form the guiding condition [$p$; $v$]. This combined condition is employed to compute the LDM’s loss function. Our optimization goal to learn the distribution parameter $v_{k,i}$ of $i_{th}$ base distribution for client $k$ is formulated as:
@@
where $z$ is the latent code of the input image $B_{k,i}$ generated by the encoder $E$, $\epsilon$ refers to noise sampled from the standard normal distribution $\mathcal{N}(0,1)$, $t$ denotes the timestep in the diffusion process, $\alpha_{t}$ is a hyperparameter related to $t$ and $\epsilon_{\theta}$ is the model that predicts the noise. Therefore, we derive the $m_{k}$ data representations $v_{k,i}$ by solving the optimization problem in Eq. (12).
@@
6:     $z_{k,i}=E(x_{k,i})$;
7:     Clustering $\mathcal{D}_{k}$ into $m_{k}$ datasets $\{B_{k,i}\}_{i=1}^{m_{k}}\}$
8:       according to $z_{k,i}$ as Eq. (11);
9:     Initialize $\{v_{k,i}\}_{i=1}^{m_{k}}$;
10:     for $i$ in [$m_{k}$do:
11:         Learn the base distribution parameter $v_{k,i}$
12:             by optimizing Eq. (12);
13:     Upload $\{\widetilde{v}_{k,i}\}_{i=1}^{m_{k}}$ to the server:
14:         $\widetilde{v}_{k,i}=v_{k,i}/max\{1,\frac{\|v_{k,i}\|_{2}}{C}\}+\mathcal{N}(0,\sigma^{2}C^{2}I)$.
15$\triangleright$ The server performs:
16:Build the binary assignment $e_{i,j}^{k_{1},k_{2}}$ via Eq. (14);
17:Obtain multiple parallel set:
18:   $\mathcal{I}_{p}=\{(k,i)|e^{k_{1},k_{2}}_{i_{1},i_{2}}=1,k\in[K],i\in[m_{k}]\}$;
19:Obtain the orthogonal set:
20:    $\mathcal{I}_{o}=\{(k,i)|k\in[K],i\in[m_{k}]\}-\mathcal{I}_{p}$
21:Obtain $v_{o}=\bigcup_{(k,i)\in\mathcal{I}_{o}}v_{k,i}$;
22:Select the distribution parameter $v_{p}$ in $\mathcal{I}_{p}$ with the
23:  maximum trained data number;
24:Distribute $v=[v_{p},v_{o}]$ to all clients.
25$\triangleright$ Clients perform:
26:for Client $k$ in $\{1,\dots,K\}$ do:
27:     Generate data $\{\widehat{x}_{i,j}|j\in[n_{i}]\}_{i=1}^{|v_{0}|}$ using $v$:
28:        $\widehat{x}_{i,j}=D(f(z_{T},T,[p,v_{o}^{i}])),j\in[n_{i}]$;
29:     Train local model using generated Data:
30:        $\mathcal{F}_{k}^{*}=\text{argmin}_{\mathcal{F}_{k}}\mathbf{E}_{i\in[|v_{0}|]}\mathcal{L}_{CE}(\mathcal{F}_{k}(\widehat{x}_{i,j}),\widehat{y}_{i})$.
31:return Downstream model $\{\mathcal{F}_{k}^{*}\}_{k=1}^{K}$.
These representations are then utilized to the diffusion model to generate the corresponding datasets for each disentangled base distribution $i$. In the federated learning setting, each client uploads their noised respective set of representations $\{\widehat{v}_{k,i}\}_{i=1}^{m_{k}}$ and corresponding data number to the server.
$\begin{split}\widetilde{v}_{k,i}=&\ v_{k,i}/max\{1,\frac{\|v_{k,i}\|_{2}}{C}\}\\ &+\mathcal{N}(0,\sigma^{2}C^{2}I),\\ &\ \ i\in[m_{k}],\text{ for }k\in[K],\end{split}$ (13)
where $C$ is the norm upper bound, $\sigma$ is the standard derivation of added noise.
4.3 Active Distribution Alignment
Moreover, since there are no explicit labels for the uploaded base distribution embeddings, matching these parameters across different clients presents a challenge. This matching problem can be framed as an assignment problem in a bipartite graph Dulmage & Mendelsohn (1958). To address this assignment problem, we utilize the Kuhn-Munkres algorithm (KM) Zhu et al. (2011), which is designed to find the maximum-weight matching in a weighted bipartite graph, or equivalently, to minimize the assignment cost. Specifically, the goal is to optimize the following expression for matching prompt embeddings between client $1$ and $k$:
$\min\sum_{i_{1}=1}^{m_{k_{1}}}\sum_{i_{2}=1}^{m_{k_{2}}}\|v_{k_{1},i_{1}}-v_{k_{2},i_{2}}\|e^{k_{1},k_{2}}_{i_{1},i_{2}},$ (14)
4.4 Local Data Generation and Downstream Training
Each client receives the distributed base distributions $v=[v_{p},v_{o}]$ from the server. For each base distribution embedding $v_{o}^{i}\in v_{o}$, clients use the latent diffusion model to generate data $\widehat{x}_{i}=\{\widehat{x}_{i,j}|j\in[n_{i}]\}$ following the base distribution.
$\widehat{x}_{i,j}=D(f(z_{T},T,[p,v_{o}^{i}])),j\in[n_{i}]\text{, for }i\in[|v_{0}|],$ (15)
where decoder $D$ is for the inverse mapping of encoder $E$, $f$ is the denoising process, $T$ is the number of diffusion steps, and $[p,v_{i}]$ is the conditional prompt corresponding to base distribution $i$.
Using the synthetic data, the clients locally train their downstream task model.
$\mathcal{F}_{k}^{*}=\text{argmin}_{\mathcal{F}_{k}}\mathbf{E}_{i\in[|v_{0}|]}\mathcal{L}_{CE}(\mathcal{F}_{k}(\widehat{x}_{i,j}),\widehat{y}_{i})\text{, for }k\in[K],$ (16)
where $\mathcal{F}_{k}$ is the local downstream model of client $k$, $\mathcal{L}_{CE}$ is the cross entropy loss, $\{\widehat{x}_{i,j}|j\in[n_{i}]\}$ is the generated data corresponding to base distribution $i$.